# InstructRetro: Instruction Tuning post Retrieval-Augmented Pretraining

### v2

@@ Abstract
Pretraining auto-regressive large language models (LLMs) with retrieval demonstrates better perplexity and factual accuracy by leveraging external databases. However, the size of existing pretrained retrieval-augmented LLM is still limited (e.g., Retro has 7.5B parameters), which limits the effectiveness of instruction tuning and zero-shot generalization. In this work, we introduce Retro 48B, the largest LLM pretrained with retrieval. Specifically, we continue to pretrain a 43B GPT model on additional 100 billion tokens using the Retro augmentation method by retrieving from 1.2 trillion tokens. Notably, the obtained foundation model, Retro 48B, largely outperforms the counterpart GPT 43B trained on 1.2T tokens in terms of perplexity with only 2.58% additional GPU hours, demonstrating the significant scaling potential of the method. After instruction tuning on Retro, InstructRetro demonstrates significant improvement over the instruction tuned GPT on a wide range of zero-shot tasks. Specifically, the average improvement of InstructRetro is 7% over its GPT counterpart across 8 short-form QA and reading comprehension tasks, 10% over GPT across 4 challenging long-form QA tasks, and 16% over GPT across 3 summarization tasks. Surprisingly, we find that one can ablate the encoder from InstructRetro architecture and directly use its decoder backbone, while achieving comparable results. Our results highlight the promising direction to obtain a better GPT decoder through continued pretraining with retrieval before instruction tuning. Our code and checkpoints are publicly available at: https://github.com/NVIDIA/Megatron-LM/tree/InstructRetro/tools/retro.
@@ 2 Related Work
Retrieval-augmented language models have been established for open domain question answering for years (Karpukhin et al., 2020; Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022b). In the previous study, language models have been augmented with retrieval at inference (Khandelwal et al., 2020; Yogatama et al., 2021), fine-tuning (Karpukhin et al., 2020; Lewis et al., 2020; Guu et al., 2020; Huang et al., 2023), and pretraining (Borgeaud et al., 2022; Izacard et al., 2022b; Wang et al., 2023a). Retrieval-augmented pretraining is particularly interesting, as it can largely reduce model perplexity (Borgeaud et al., 2022), enhance factuality (Wang et al., 2023a), and improve downstream task accuracy after task-specific fine-tuning (Izacard et al., 2022b).
@@
Instruction tuning aims to teach LLMs to follow natural language instructions (Wei et al., 2022a; Ouyang et al., 2022; Sanh et al., 2022b; Mishra et al., 2022), which becomes an indispensable ingredient to build the state-of-the-art LLMs for downstream tasks (OpenAI, 2022, 2023; Touvron et al., 2023b). In the past years, many high-quality instruction tuning datasets have been created, including FLAN (Chung et al., 2022), OpenAssistant (Köpf et al., 2023), Self-Instruct (Wang et al., 2022a), Dolly (Conover et al., 2023), Unnatural Instructions (Honovich et al., 2022).
@@ 5.1 Experimental setup
Datasets. To demonstrate the generalization of instruction tuning, we follow FLAN (Wei et al., 2022a) and primarily focus on zero-shot evaluation of downstream tasks. Specifically, we consider two categories of open-ended QA tasks as well as text summarization tasks: (1) short-form QA or reading comprehension, which expects short answers (e.g., a few tokens) to be generated or extracted from the context, including Natural Question (NQ) (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), NewsQA (Trischler et al., 2016), SQuAD 1.1 (Rajpurkar et al., 2016), SQuAD 2.0 (Rajpurkar et al., 2018), Quoref (Dasigi et al., 2019), NarrativeQA (Kočiskỳ et al., 2018), DROP (Dua et al., 2019); (2) long-form QA, which expects longer answer spans within a few sentences, including doc2dial (Feng et al., 2020), two proprietary annotated car manual datasets (people ask questions about the particular car models), and another proprietary annotated IT documentation dataset; (3) summarization, which expects to summarize a long passage or context within a few sentences, including QMSum (Zhong et al., 2021), SummScreenFd (Chen et al., 2021), and GovReport (Huang et al., 2021).
Retrieval-augmented generations (RAG). At pretraining, we use BERT embeddings to embed the retrieval database and support retrieval from trillions of tokens. For downstream task evaluation, we follow Retro (Borgeaud et al., 2022) and use task-specific corpus and state-of-the-art retrievers (e.g., DPR (Karpukhin et al., 2020) and DRAGON+ (Lin et al., 2023)) to retrieve the most relevant and high-quality information for the task. For NQ and TriviaQA, we use DPR (Karpukhin et al., 2020) to retrieve from Wikipedia. For doc2dial and other long-form QA datasets, we use DRAGON+ (Lin et al., 2023). We retrieve the top- nearest neighbors and concatenate them in the prompt. For the remaining QA and summarization tasks, we use the provided contexts in the datasets. An example of how we format the retrieved neighbors in the prompt is shown in Appendix Table 10.
@@ 5.4 Ablation stuides
To remove the instruction-following bottleneck, we apply instruction tuning to further fine-tune both Retro 48B and GPT 43B. Instruction tuning largely mitigate the instruction following bottleneck for both GPT and Retro, resulting in a significant increase of their zero-shot performance on downstream tasks, respectively. Furthermore, once this bottleneck is alleviated, the benefits of retrieval augmentation at pretraining become more pronounced, as InstructRetro excels in leveraging and integrating evidence from retrieved context. Thus, we observe significant improvement of InstructRetro over Instruct again in Figure 4(b). This ablation study confirms that our training recipe - both retrieval-augmented pretraining and instruction tuning are important for achieving high performance in QA tasks. We also conduct additional ablation studies on the impact of retrieval-augmentation generation in Appendix §D.1.
@@ C.1 An example From the instruction tuning data
System |
This is a chat between a user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions based on the context. The assistant should also indicate when the answer cannot be found in the context. |
User |
My bicycle tire has a puncture and while I have a pump with me, I don’t have a patch repair kit. Is there a way to keep the tire from running out of air until I can get to a bicycle shop? |
Assistant |
Yes, you could use duct tape to seal the puncture. |
User |
Can you give me a tutorial on how to use the duct tape to seal the puncture? |
@@
System |
System: This is a chat between a user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions based on the context. The assistant should also indicate when the answer cannot be found in the context. |
@@
Assistant |
The answer is |
Appendix D Additional experimental results
D.1 Ablation studies on retrieval augmentation
We apply retrieval-augmented generation (RAG) across our main experiments. In this subsection, we conduct ablation studies on the impact of RAG on open-domain QA datasets Natural Question (NQ) and TriviaQA for both GPT-Instruct and InstructRetro. We set the Retro encoder gate OFF in the ablation studies. The results are shown in Table 12 and 12.
| EM Score | GPT-Instruct | InstructRetro |
| w/o RAG | 21.2 | 21.8 |
| w/ RAG | 37.0 | 38.9 |
| EM Score | GPT-Instruct | InstructRetro |
| w/o RAG | 53.6 | 54.5 |
| w/ RAG | 65.0 | 65.6 |
From Table 12 and 12, we can see that the accuracy gap between with or without RAG is significant. For example, the EM scores of NQ surges from 21.8 to 38.9 for InstructRetro. The improvement margin is significant GPT-Instruct as well. Moreover, we observe that InstructRetro consistently outperforms GPT-Instruct across different tasks, even without retrieval augmentation. This further confirms that instruction tuning after retrieval-augmented pretraining can help yield a better GPT decoder.

### v3

@@ Abstract
Pretraining auto-regressive large language models (LLMs) with retrieval demonstrates better perplexity and factual accuracy by leveraging external databases. However, the size of existing pretrained retrieval-augmented LLM is still limited (e.g., Retro has 7.5B parameters), which limits the effectiveness of instruction tuning and zero-shot generalization. In this work, we introduce Retro 48B, the largest LLM pretrained with retrieval. Specifically, we continue to pretrain a 43B GPT model on additional 100 billion tokens using the Retro augmentation method by retrieving from 1.2 trillion tokens. Notably, the obtained foundation model, Retro 48B, largely outperforms the counterpart GPT 43B trained on 1.2T tokens in terms of perplexity with only 2.58% additional GPU hours, demonstrating the significant scaling potential of the method. After instruction tuning on Retro, InstructRetro demonstrates significant improvement over the instruction tuned GPT on a wide range of zero-shot tasks. Specifically, the average improvement of InstructRetro is 7% over its GPT counterpart across 8 short-form QA and reading comprehension tasks, 10% over GPT across 4 challenging long-form QA tasks, and 16% over GPT across 3 summarization tasks. Surprisingly, we find that one can ablate the encoder from InstructRetro architecture and directly use its decoder backbone, while achieving comparable results. Our results highlight the promising direction to obtain a better GPT decoder through continued pretraining with retrieval before instruction tuning. Our code and checkpoints are publicly available at: https://huggingface.co/nvidia/retro-48b-instruct-4k.
@@ 2 Related Work
Retrieval-augmented language models have been established for open domain question answering for years (Karpukhin et al., 2020; Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022b). In the previous study, language models have been augmented with retrieval at inference (Khandelwal et al., 2020; Yogatama et al., 2021), fine-tuning (Karpukhin et al., 2020; Lewis et al., 2020; Guu et al., 2020; Huang et al., 2023; Shi et al., 2023b), and pretraining (Borgeaud et al., 2022; Izacard et al., 2022b; Wang et al., 2023a; Shi et al., 2023a). Retrieval-augmented pretraining is particularly interesting, as it can largely reduce model perplexity (Borgeaud et al., 2022), enhance factuality (Wang et al., 2023a), and improve downstream task accuracy after task-specific fine-tuning (Izacard et al., 2022b) and reasoning capability (Shi et al., 2023a).
@@
Instruction tuning aims to teach LLMs to follow natural language instructions (Wei et al., 2022a; Ouyang et al., 2022; Sanh et al., 2022b; Mishra et al., 2022), which becomes an indispensable ingredient to build the state-of-the-art LLMs for downstream tasks (OpenAI, 2022, 2023; Touvron et al., 2023b). In the past years, many high-quality instruction tuning datasets have been created, including FLAN (Chung et al., 2022), OpenAssistant (Köpf et al., 2023), Self-Instruct (Wang et al., 2022a), Dolly (Conover et al., 2023), Unnatural Instructions (Honovich et al., 2022). A concurrent work, RA-DIT (Lin et al., 2024), focuses on retrieval-augmented instruction tuning and further augments 20 instruction tuning datasets with retrieval, which supports fine-tuning both LLM and retriever to yield high-quality neighbors. In contrast, our work focuses on retrieval-augmented pretraining, which extends the scale of the retrieval database to trillions of tokens. Although the two work are orthogonal, InstructRetro 43B outperforms RA-DIT 65B on certain benchmarks as shown in Table 1. We leave it as an interesting future direction to apply RA-DIT retrieval-augmented instruction tuning data to the instruction tuning stage of InstructRetro for further performance improvement of retrieval-augmented LLMs.
@@ 5.1 Experimental setup
Datasets. To demonstrate the generalization of instruction tuning, we follow FLAN (Wei et al., 2022a) and primarily focus on zero-shot evaluation of downstream tasks. Specifically, we consider two categories of open-ended QA tasks as well as text summarization tasks: (1) short-form QA or reading comprehension, which expects short answers (e.g., a few tokens) to be generated or extracted from the context, including Natural Question (NQ) (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), NewsQA (Trischler et al., 2016), SQuAD 1.1 (Rajpurkar et al., 2016), SQuAD 2.0 (Rajpurkar et al., 2018), Quoref (Dasigi et al., 2019), NarrativeQA (Kočiskỳ et al., 2018), DROP (Dua et al., 2019). To compare with baselines, we use the split from KILT benchmark (Petroni et al., 2021) for NQ and TriviaQA. For the other tasks, we use the official splits; (2) long-form QA, which expects longer answer spans within a few sentences, including doc2dial (Feng et al., 2020), two proprietary annotated car manual datasets (people ask questions about the particular car models), and another proprietary annotated IT documentation dataset; (3) summarization, which expects to summarize a long passage or context within a few sentences, including QMSum (Zhong et al., 2021), SummScreenFd (Chen et al., 2021), and GovReport (Huang et al., 2021).
Retrieval-augmented generations (RAG). At pretraining, we use BERT embeddings to embed the retrieval database and support retrieval from trillions of tokens. For downstream task evaluation, we follow Retro (Borgeaud et al., 2022) and use task-specific corpus and state-of-the-art retrievers to retrieve the most relevant and high-quality information for the task. Specifically, for NQ, TriviaQA, doc2dial, and other long-form QA datasets, we use DRAGON+ (Lin et al., 2023) as the retriever. We retrieve the top- nearest neighbors and concatenate them in the prompt. For the remaining QA and summarization tasks, we use the provided contexts in the datasets. An example of how we format the retrieved neighbors in the prompt is shown in Appendix Table 10.
@@ 5.4 Ablation stuides
To remove the instruction-following bottleneck, we apply instruction tuning to further fine-tune both Retro 48B and GPT 43B. Instruction tuning largely mitigate the instruction following bottleneck for both GPT and Retro, resulting in a significant increase of their zero-shot performance on downstream tasks, respectively. Furthermore, once this bottleneck is alleviated, the benefits of retrieval augmentation at pretraining become more pronounced, as InstructRetro excels in leveraging and integrating evidence from retrieved context. Thus, we observe significant improvement of InstructRetro over Instruct again in Figure 5(b). This ablation study confirms that our training recipe - both retrieval-augmented pretraining and instruction tuning are important for achieving high performance in QA tasks.
5.5 Evaluation on MT-Bench
@@ C.1 An example From the instruction tuning data
| System | This is a chat between a user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions based on the context. The assistant should also indicate when the answer cannot be found in the context. |
| User | My bicycle tire has a puncture and while I have a pump with me, I don’t have a patch repair kit. Is there a way to keep the tire from running out of air until I can get to a bicycle shop? |
| Assistant | Yes, you could use duct tape to seal the puncture. |
| User | Can you give me a tutorial on how to use the duct tape to seal the puncture? |
@@
1. Start by cleaning the area around the puncture with a damp cloth. 2. Cut a piece of duct tape large enough to cover the puncture, plus a few millimeters of extra. 3. Remove the backing from the duct tape, and press it firmly over the puncture. 4. Rub the tape firmly with your fingers to ensure a good seal. 5. You can also use a roller or a flat item like a ruler to make sure that the tape is pressed firmly against the puncture. 6. Trim away any excess tape, and your puncture should now be sealed. |
@@
| System | System: This is a chat between a user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions based on the context. The assistant should also indicate when the answer cannot be found in the context. |
@@
| Assistant | The answer is |
Appendix D Experimental results on MT Bench
We evaluate InstructRetro and InstructGPT on the MTBench chat benchmark (Zheng et al., 2024) to understand Retro performance on general chat tasks. We use the Tulu-v2 (Wang et al., 2023b) alignment dataset as the instruction tuning dataset to enhance the chat capabilities for both pretrained GPT model and Retro model. The detailed breakdown of the MT-Bench result is shown in Table 11.
| MT-Bench | InstructRetro-Tulu-v2-43B | InstructGPT-Tulu-v2-43B |
| Writing | 8.85 | 8.15 |
| Roleplay | 7.75 | 7.80 |
| Reasoning | 5.40 | 4.75 |
| Math | 3.15 | 2.35 |
| Coding | 3.40 | 4.10 |
| Extraction | 6.80 | 6.75 |
| STEM | 8.58 | 8.53 |
| Humanities | 9.68 | 9.10 |
| Turn 1 | 6.89 | 6.67 |
| Turn 2 | 6.51 | 6.21 |
| Avg | 6.70 | 6.44 |
In Table 11, we show that on average InstructRetro outperforms InstructGPT across different turns. We also observe that Retro performs slightly lower than GPT in domains such as role play and coding. We think that the main reason is that we do not cover coding and role-playing related datasets in the pretraining dataset of the base GPT model, and thus retrieval-augmented pretraining could make little difference.